# Tropicalizing Binary Geometries

### v1

@@ Introduction
The notion of binary geometries was introduced in [3] in the context of particle scattering in arbitrary space-time dimension. In particle physics it arises from the ABHY kinematic associahedron ([1]) and its extension [4] to generalized associahedra of all finite-type cluster algebras (in the sense of [8, 7]).
@@
From a purely cluster-algebraic standpoint one may view binary geometries as partial compactifications of torus quotients of cluster varieties, i.e. cluster configuration spaces, cf. also [11, 2]. They are realized as affine varieties cut out by the so-termed $u$-equations, which are explicit defining relations determined by the combinatorics of the generalized associahedron. Importantly, the resulting variety admits a stratification governed by the face poset of the same polytope. In [12] this definition is widely generalized to the setting of flag simplicial complexes. In this paper we choose the latter approach, although our main objects of study are the binary geometries given by the type A and type C associahedra and the respective finite-type cluster configuration spaces.
@@
Let us first consider type A. The cluster configuration space $\mathcal{M}_{\mathrm{A}_{n-3}}$ is usually denoted by $\mathcal{M}_{0,n}$ and is the moduli space of $n$-tuples of distinct points in $\mathbb{P}^{1}$. We view it as the very affine part of the binary geometry ~$\widetilde{\mathcal{M}}_{0,n}$ associated with the complex dual to the associahedron. The tropicalization of $\mathcal{M}_{0,n}$ (and, thus, of ~$\widetilde{\mathcal{M}}_{0,n}$) is well-studied; its description can be found in [9] and the book [13, Section 6.4] but traces back at least to [14, 16]. This description characterizes it as the space of phylogenetic trees introduced in [5]. Furthermore, this space is naturally the union of $(n-1)!/2$ dual-associahedron subfans. These are precisely the positive subfans (or signed tropicalizations) corresponding to all sign patterns of coordinates occurring in $\mathcal{M}_{0,n}(\mathbb{R})$. In the first sections of this paper we recall these notions and results and write explicit formulas for the cones of $\operatorname{Trop}\mathcal{M}_{0,n}$ in the $u$-coordinates arising from the binary geometry structure. This overview provides a background for our main results which concern type C.
@@
Our first main result (see Sections 56) is the construction of an adequate analogue of the space of phylogenetic trees for the symplectic setting. The type A space is the fan over a simplicial complex with faces enumerated by phylogenetic trees with $n$ labeled leaves. In type C we instead consider trees with $2n$ labeled leaves that are axially symmetric in a natural sense. We define the complex of axially symmetric phylogenetic trees $\Theta_{\mathrm{as}}(n)$, a simplicial complex of dimension $n-2$ with faces enumerated by such trees. It is naturally the union of $2^{n-2}n!$ dual-associahedron subcomplexes. We also consider trees with $2n$ labeled leaves that are centrally symmetric. The resulting complex of centrally symmetric phylogenetic trees $\Theta_{\mathrm{cs}}(n)$ is shown to be a subcomplex of $\Theta_{\mathrm{as}}(n)$. Furthermore, the subcomplex $\Theta_{\mathrm{cs}}(n)$ is itself the union of $2^{n-2}(n-1)!$ dual-cyclohedron subcomplexes.
@@
Now, the cyclohedron is the generalized associahedron of type C. Accordingly, $\mathcal{M}_{\mathrm{C}_{n-1}}$ is the very affine part of the binary geometry defined by the cyclohedron’s dual complex (the shift of numeration is needed to match the conventions in the previous paragraph). This binary geometry is an affine variety of dimension $n-1$. Our second main result is a conjecture.
@@
The tropicalization of $\mathcal{M}_{\mathrm{C}_{n-1}}$ is combinatorially a fan over the complex of axially symmetric phylogenetic trees $\Theta_{\mathrm{as}}(n)$. Furthermore the subfans given by the $2^{n-2}n!$ dual associahedra and the $2^{n-2}(n-1)!$ dual cyclohedra are the signed tropicalizations corresponding to all sign patterns of coordinates occurring in $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$.
The main motivation for the conjecture is the case $n=3$ which is already highly nontrivial: the tropicalization has 21 maximal cones and 13 rays (as found using Macaulay2, [10]). Moreover, in agreement with this computation we conjecture an explicit $u$-coordinate description of the cone corresponding to every symmetric tree.
A proof of our conjecture would help uncover further geometric structure in $\mathcal{M}_{\mathrm{C}_{n-1}}$. In particular, in [2] it is conjectured that the occurring sign patterns are in bijection with the connected components of $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$ and that their number agrees with that provided by Conjecture 0.1.
@@ 1 Tropicalizations
Note that the tropicalizations of $\varepsilon_{\tau}(I)$ and $I$ are the same, hence $\operatorname{Trop}_{\tau}I$ is also a subfan of $\operatorname{Trop}I$. Furthermore, one sees that $\operatorname{Trop}_{\tau}I$ is nonempty whenever the sign pattern $\tau$ occurs in $X(\mathbb{R})=X\cap\mathbb{R}^{V}$, which means that there exists a point $x\in X(\mathbb{R})$ such that $\operatorname{sgn}x_{i}=\tau_{i}$ for all $i$.
@@ 2 Phylogenetic trees and dual associahedra
Choose an integer $n\geq 3$.
@@
Choose a dihedral ordering $\alpha$, and consider a convex $n$-gon $P$ with edges labeled in accordance with $\alpha$. Every subdivision of $P$ defines a phylogenetic tree as follows. First consider the adjacency graph of the 2-dimensional cells of the subdivision, this is a tree. Then, for the cell $C$ containing the edge of $P$ labeled by $i\in[n]$ add a leaf adjacent to the vertex corresponding to $C$, label this leaf by $i$ (see figure below). The result is a phylogenetic tree with $n$ leaves. We say that a phylogenetic tree is compatible with $\alpha$ if it arises from a subdivision of $P$ in the above way.
111122223333444455556666
@@
We see that the faces of $\Delta(\alpha)$ are in bijection with all subdivisions of $P$ and, therefore, with all phylogenetic trees compatible with $\alpha$111 We consider the empty set to be a face of any simplicial complex, here it corresponds to the trivial subdivision and the phylogenetic tree with a single non-leaf vertex. This is a flag complex of pure dimension $n-4$ (recall that a simplicial complex is a flag complex if all of its minimal non-faces are edges).222 For $n=3$ the complex $\Delta(\alpha)$ is empty and thus has a single face and is of dimension $-1$. We refer to $\Delta(\alpha)$ as the dual associahedron because its face poset is dual to that of an associahedron.
We denote $\Delta(n)=\Delta(\alpha)$ for $\alpha$ the dihedral ordering given by the identity permutation. Obviously, for a fixed $n$ every $\Delta(\alpha)$ is isomorphic to $\Delta(n)$. However, we view it as a distinct complex because it has a distinct vertex set (though not necessarily disjoint since a phylogenetic tree will be compatible with multiple dihedral orderings). In particular, this allows for the following definition.
@@
The complex of phylogenetic trees $\Theta(n)$ is the union of $\Delta(\alpha)$ over all $\alpha$. This is the simplicial complex whose vertex set consists of all phylogenetic trees with $n$ leaves and exactly two non-leaf vertices. A subset of vertices forms a face if and only if it forms a face in one of the $\Delta(\alpha)$.
The faces of $\Theta(n)$ are in bijection with all phylogenetic trees with $n$ leaves which explains the terminology. Evidently, $\Theta(n)$ is also a flag complex of pure dimension $n-4$. The fan over this simplicial complex is known as the space of phylogenetic trees, a term introduced in [5] (occasionally $\Theta(n)$ is referred to by the same name but we distinguish between the two).
@@ 3 Binary geometries
Note that $\widetilde{U}_{\varnothing}=\widetilde{U}$, hence, $\widetilde{U}$ is irreducible of dimension $\dim\Delta+1$. Somewhat less trivially, one may show that if $F$ is a facet then $\widetilde{U}_{F}$ is a point. This follows from the fact that if $\widetilde{U}$ is a binary geometry, then $\Delta$ is pure ([12, Corollary 2.7]). Indeed, for every facet $F$ we have $|F|=\dim\Delta+1$ and consequently $\dim U_{F}=0$.
@@ 4 $\mathcal{M}_{0,n}$ and its tropicalization
Definition 4.1 of $\mathcal{M}_{0,n}$ is consistent with the definition of $\mathcal{M}_{0,n}$ as a quotient of $\operatorname{Gr}^{\circ}(2,n)$, see [6]. Given a point in $\mathcal{M}_{0,n}$, i.e. an $n$-tuple $(x_{1},\ldots,x_{n})$ of points in $\mathbb{P}^{1}$, the coordinate $u_{i,j}$ is equal to the cross-ratio $(x_{i},x_{i+1};x_{j},x_{j+1})$. Therefore, the projection from $\operatorname{Gr}^{\circ}(2,n)$ to $\mathcal{M}_{0,n}$ is given by the formula
@@
We now turn to the discussion of sign patterns and signed tropicalizations of $\mathcal{M}_{0,n}$. First of all, note that an element of the symmetric group $\mathfrak{S}_{n}$ acts naturally on $\mathcal{M}_{0,n}$ by permuting the points of the respective $n$-tuple in $\mathbb{P}^{1}$ (i.e. the columns of the $2\times n$ matrix). With respect to the coordinates $u_{i,j}$ every permutation acts by a monomial transformation, see [2, Subsection 11.2]. This induces an $\mathfrak{S}_{n}$-action on $\operatorname{Trop}\mathcal{M}_{0,n}$ by linear transformations of the ambient space $\mathbb{R}^{D}$. Furthermore, $\mathfrak{S}_{n}$ also acts on the set of phylogenetic trees by $\sigma(T,\varphi)=(T,\varphi\circ\sigma^{-1})$ for $\sigma\in\mathfrak{S}_{n}$. By tracing the definitions one sees that this action is compatible with the action on $\operatorname{Trop}\mathcal{M}_{0,n}$, i.e. $\sigma(C_{T,\varphi})=C_{\sigma(T,\varphi)}$.
There are $(n-1)!/2$ sign patterns occurring in $\mathcal{M}_{0,n}(\mathbb{R})$, which are enumerated by dihedral orderings of $[n]$, again, see [2, Subsection 11.2]. For a dihedral ordering $\alpha$ we denote the corresponding sign pattern by $\tau_{\alpha}$ and the set of points in $\mathcal{M}_{0,n}$ with coordinates fitting this sign pattern by $\mathcal{M}_{0,n}(\alpha)$. In fact, the subsets $\mathcal{M}_{0,n}(\alpha)$ are precisely the connected components of $\mathcal{M}_{0,n}(\mathbb{R})$. The natural action of $\mathfrak{S}_{n}$ on the set of dihedral orderings is compatible with its action on $\mathcal{M}_{0,n}$: one has $\sigma(\mathcal{M}_{0,n}(\alpha))=\mathcal{M}_{0,n}(\sigma\alpha)$ so that the action permutes the connected components transitively. This means that the corresponding signed tropicalizations given by the $\tau_{\alpha}$ are also permuted by the $\mathfrak{S}_{n}$-action: $\operatorname{Trop}_{\tau_{\sigma\alpha}}\mathcal{M}_{0,n}=\sigma(\operatorname{Trop}_{\tau_{\alpha}}\mathcal{M}_{0,n})$.

### v2

@@ Introduction
The notion of binary geometries was introduced in [3] in the context of particle scattering in arbitrary space-time dimension. In particle physics it arises from the ABHY kinematic associahedron ([1]) and its extension [4] to generalized associahedra of other finite-type cluster algebras (in the sense of [8, 7]).
@@
From a cluster-algebraic standpoint one may view binary geometries as cluster configuration spaces, i.e. partial compactifications of torus quotients of cluster varieties, cf. also [11, 2]. Such binary geometries are realized as affine varieties cut out by the so-termed $u$-equations: defining relations explicitly determined by the combinatorics of the generalized associahedron. In particular, the resulting variety admits a stratification governed by the face poset of the same polytope. In [12] this definition is widely generalized to the setting of flag simplicial complexes. In this paper we follow the approach in [12], nonetheless, our main objects of study are the binary geometries given by the type A and type C associahedra, and the respective finite-type cluster configuration spaces.
@@
Let us first consider type A. The cluster configuration space $\mathcal{M}_{\mathrm{A}_{n-3}}$ is commonly denoted by $\mathcal{M}_{0,n}$; it is the moduli space of $n$-tuples of distinct points in $\mathbb{P}^{1}$. It may also be viewed as the very affine part of the binary geometry ~$\widetilde{\mathcal{M}}_{0,n}$ associated with the simplial complex dual to the associahedron. The tropicalization of $\mathcal{M}_{0,n}$ (and, thus, of ~$\widetilde{\mathcal{M}}_{0,n}$) is well-studied; its description can be found in [9], the book [13, Section 6.4] and other sources but traces back at least to [15, 17]. This description characterizes $\operatorname{Trop}\mathcal{M}_{0,n}$ as the space of phylogenetic trees introduced in [14, 5]. Furthermore, $\operatorname{Trop}\mathcal{M}_{0,n}$ is naturally realized as the union of $(n-1)!/2$ dual-associahedron subfans. These are precisely the signed tropicalizations (i.e. generalized positive tropicalizations) corresponding to all sign patterns of coordinates occurring in $\mathcal{M}_{0,n}(\mathbb{R})$. In the first sections of this paper we recall the definitions and results pertaining to $\mathcal{M}_{0,n}$ and its tropicalization. We also write down explicit formulas for the cones of $\operatorname{Trop}\mathcal{M}_{0,n}$ in the $u$-coordinates arising from the binary geometry structure. This overview provides a background and motivation for our main results which concern type C.
@@
Our first result (see Sections 56) is the construction of an adequate analogue of the space of phylogenetic trees for the symplectic setting. The previously mentioned description of $\operatorname{Trop}\mathcal{M}_{0,n}$ realizes it as the fan over a simplicial complex with faces enumerated by phylogenetic trees with $n$ labeled leaves. In type C we, instead, consider trees with $2n$ labeled leaves that are axially symmetric in a natural sense. We define the complex of axially symmetric phylogenetic trees $\Theta_{\mathrm{as}}(n)$, a simplicial complex of dimension $n-2$ with faces enumerated by such trees. It is naturally the union of $2^{n-2}n!$ dual-associahedron subcomplexes. We also consider trees with $2n$ labeled leaves that are centrally symmetric. The resulting complex of centrally symmetric phylogenetic trees $\Theta_{\mathrm{cs}}(n)$ is shown to be a subcomplex of $\Theta_{\mathrm{as}}(n)$. Furthermore, the subcomplex $\Theta_{\mathrm{cs}}(n)$ is itself the union of $2^{n-2}(n-1)!$ dual-cyclohedron subcomplexes.
@@
Next, recall that the cyclohedron is the generalized associahedron of type $\mathrm{C}_{n-1}$. Accordingly, the cluster configuration space $\mathcal{M}_{\mathrm{C}_{n-1}}$ can be realized in $u$-coordinates as the very affine part of the binary geometry defined by the dual-cyclohedron complex, see [2]. This binary geometry is an affine variety of dimension $n-1$. Our second result is a conjectural description of its tropicalization.
@@
With respect to the $u$-coordinates, $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}$ is combinatorially equivalent to a fan over the complex of axially symmetric phylogenetic trees $\Theta_{\mathrm{as}}(n)$. Furthermore, the subfans given by the $2^{n-2}n!$ dual associahedra and the $2^{n-2}(n-1)!$ dual cyclohedra are the signed tropicalizations corresponding to all sign patterns of coordinates occurring in $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$.
In particular, the cones of $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}$ correspond to faces of $\Theta_{\mathrm{as}}(n)$. We also conjecture an explicit description in $u$-coordinates for every such cone. This proposed description should be viewed as an extension of the $u$-coordinate realization that was given for $\operatorname{Trop}\mathcal{M}_{0,n}$.
The main motivation for our conjectures is the case $n=3$ which is already highly nontrivial: $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{2}}$ is a fan in ambient dimension 6 with 21 maximal cones and 13 rays. This tropicalization can be found using Macaulay2 ([10]) and agrees fully with the hypothesized structure. In addition, the number of sign patterns provided by Conjecture 0.1 agrees with a conjecture made in [2].
We hope the complex $\Theta_{\mathrm{as}}(n)$ and our conjecture, if proved, to have several applications. For example, it is natural to expect $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}$ to be directly related to the type C tropicalized cluster variety. Also, the well-studied Deligne–Mumford compactification ¯$\overline{\mathcal{M}}_{0,n}$ has a stratification governed by the face poset of $\operatorname{Trop}\mathcal{M}_{0,n}$. One may hope for a similar construction to exist in type C.
@@ 1 Tropicalizations
Note that the tropicalizations of $\varepsilon_{\tau}(I)$ and $I$ are the same, hence $\operatorname{Trop}_{\tau}I$ is also a subfan of $\operatorname{Trop}I$. Furthermore, one sees that $\operatorname{Trop}_{\tau}I$ is nonempty if the sign pattern $\tau$ occurs in $X(\mathbb{R})=X\cap\mathbb{R}^{V}$, which means that there exists a point $x\in X(\mathbb{R})$ such that $\operatorname{sgn}x_{i}=\tau_{i}$ for all $i$.
@@ 2 Phylogenetic trees and dual associahedra
We overview the definition and basic properties of the complex (or space) of phylogenetic trees constructed in [14, 5]. Choose an integer $n\geq 3$.
@@
Choose a dihedral ordering $\alpha$, and consider a convex $n$-gon $P$ with edges labeled in accordance with $\alpha$. Every subdivision of $P$ defines a phylogenetic tree as follows. First consider the adjacency graph of the 2-dimensional cells of the subdivision, this is a tree. Then, for the cell $C$ containing the edge of $P$ labeled by $i\in[n]$ add a leaf adjacent to the vertex corresponding to $C$, label this leaf by $i$ (see figure on the right). The result is a phylogenetic tree with $n$ leaves. We say that a phylogenetic tree is compatible with $\alpha$ if it arises from a subdivision of $P$ in the above way.
@@
We see that the faces of $\Delta(\alpha)$ are in bijection with all subdivisions of $P$ and, therefore, with all phylogenetic trees compatible with $\alpha$111 We consider the empty set to be a face of any simplicial complex, here it corresponds to the trivial subdivision and the phylogenetic tree with a single non-leaf vertex. This is a flag complex of pure dimension $n-4$ (recall that a simplicial complex is a flag complex if all of its minimal non-faces are of size two).222 For $n=3$ the complex $\Delta(\alpha)$ is empty and thus has a single face and is of dimension $-1$. We refer to $\Delta(\alpha)$ as a dual associahedron because its face poset is dual to that of an associahedron. For example, if $\alpha$ is any dihedral ordering of five elements, $\Delta(\alpha)$ is a 5-cycle.
We denote $\Delta(n)=\Delta(\alpha)$ for $\alpha$ the dihedral ordering given by the identity permutation. Obviously, for a fixed $n$ every $\Delta(\alpha)$ is isomorphic to $\Delta(n)$, however, we view the various $\Delta(\alpha)$ as distinct complexes because they have distinct vertex sets. It is important to note that these vertex sets are not necessarily disjoint since a phylogenetic tree can be realized in the plane in different ways, meaning that it will be compatible with multiple dihedral orderings. In particular, this allows for the following definition.
@@
The complex of phylogenetic trees $\Theta(n)$ is the union of $\Delta(\alpha)$ over all $\alpha$, with vertices represented by the same phylogenetic trees identified. It is the simplicial complex whose vertex set consists of all phylogenetic trees on $n$ leaves that have exactly two non-leaf vertices. A subset of vertices forms a face if and only if it forms a face in one of the $\Delta(\alpha)$.
The faces of $\Theta(n)$ are in bijection with all phylogenetic trees with $n$ leaves which explains the terminology. Evidently, $\Theta(n)$ is also a flag complex of pure dimension $n-4$. The fan over this simplicial complex is known as the space of phylogenetic trees, a term introduced in [5] (occasionally $\Theta(n)$ is referred to by the same name but we distinguish between the two). For example, $\Theta(5)$ is the Petersen graph, see [5, FIG. 13]. The Petersen graph contains 12 copies of the 5-cycle, and these are precisely the 12 subcomplexes $\Delta(\alpha)$ corresponding to the 12 dihedral orderings.
@@ 3 Binary geometries
Note that $\widetilde{U}_{\varnothing}=\widetilde{U}$, hence, $\widetilde{U}$ is irreducible of dimension $\dim\Delta+1$. Somewhat less trivially, one may show that if $F$ is a facet then $\widetilde{U}_{F}$ is a point. This follows from the fact that if $\widetilde{U}$ is a binary geometry, then $\Delta$ is pure ([12, Corollary 2.7]). Indeed, for every facet $F$ we have $|F|=\dim\Delta+1$ and, consequently, $\dim\widetilde{U}_{F}=0$.
@@ 4 $\mathcal{M}_{0,n}$ and its tropicalization
Definition 4.1 of $\mathcal{M}_{0,n}$ is consistent with the definition of $\mathcal{M}_{0,n}$ as a quotient of $\operatorname{Gr}^{\circ}(2,n)$, see [6]. Given a point in $\mathcal{M}_{0,n}$, i.e. an $n$-tuple $(x_{1},\dots,x_{n})$ of points in $\mathbb{P}^{1}$, the coordinate $u_{i,j}$ is equal to the cross-ratio $(x_{i},x_{i+1};x_{j},x_{j+1})$. Therefore, the projection from $\operatorname{Gr}^{\circ}(2,n)$ to $\mathcal{M}_{0,n}$ is given by the formula
@@
We now turn to the discussion of sign patterns and signed tropicalizations of $\mathcal{M}_{0,n}$. This discussion is kept brief since it is logically independent from our results. Theorem 4.4 is only needed as a prototype that illustrates our proposed generalization to type C.
First of all, note that an element of the symmetric group $\mathfrak{S}_{n}$ acts naturally on $\mathcal{M}_{0,n}$ by permuting the points of the respective $n$-tuple in $\mathbb{P}^{1}$ (i.e. the columns of the $2\times n$ matrix). With respect to the coordinates $u_{i,j}$ every permutation acts by a monomial transformation, see [2, Subsection 11.2]. This induces an $\mathfrak{S}_{n}$-action on $\operatorname{Trop}\mathcal{M}_{0,n}$ by linear transformations of the ambient space $\mathbb{R}^{D}$. Furthermore, $\mathfrak{S}_{n}$ also acts on the set of phylogenetic trees by $\sigma(T,\varphi)=(T,\varphi\circ\sigma^{-1})$ for $\sigma\in\mathfrak{S}_{n}$. By tracing the definitions one sees that this action is compatible with the action on $\operatorname{Trop}\mathcal{M}_{0,n}$, i.e. $\sigma(C_{T,\varphi})=C_{\sigma(T,\varphi)}$.
There are $(n-1)!/2$ sign patterns occurring in $\mathcal{M}_{0,n}(\mathbb{R})$, which are enumerated by dihedral orderings of $[n]$, again, see [2, Subsection 11.2]. For a dihedral ordering $\alpha$ we denote the corresponding sign pattern by $\tau_{\alpha}$444 The correspondence $\alpha\mapsto\tau_{\alpha}$ is not hard to describe. One considers $n$ distinct points $x_{1},\dots,x_{n}$ in the real projective line ordered according to $\alpha$ and sets $(\tau_{\alpha})_{i,j}$ equal to $\operatorname{sgn}(x_{i},x_{i+1};x_{j},x_{j+1})$. and the set of points in $\mathcal{M}_{0,n}(\mathbb{R})$ with coordinates fitting the sign pattern $\tau_{\alpha}$ by $\mathcal{M}_{0,n}(\alpha)$. In fact, the subsets $\mathcal{M}_{0,n}(\alpha)$ are precisely the connected components of $\mathcal{M}_{0,n}(\mathbb{R})$. The natural action of $\mathfrak{S}_{n}$ on the set of dihedral orderings is compatible with its action on $\mathcal{M}_{0,n}$: one has $\sigma(\mathcal{M}_{0,n}(\alpha))=\mathcal{M}_{0,n}(\sigma\alpha)$ so that the action permutes the connected components transitively. This means that the corresponding signed tropicalizations given by the $\tau_{\alpha}$ are also permuted by the $\mathfrak{S}_{n}$-action: $\operatorname{Trop}_{\tau_{\sigma\alpha}}\mathcal{M}_{0,n}=\sigma(\operatorname{Trop}_{\tau_{\alpha}}\mathcal{M}_{0,n})$.